# NaP2: A Benchmark for Naturalness and Privacy-Preserving Text Rewriting by Learning from Human

### v1

@@ 4.2 Evaluation Details
Privacy Leakage. We propose a novel metric, called Privacy_NLI, by using the RoBERTa model trained on the MNLI corpus, to infer to what degree it is possible to infer personal information in personas. As the NLI model classifies a pair of input texts into entailed, contradicted, or neutral, we adopt $P(\text{entailed}|{\bm{x}},{\bm{p}})$ as the score of $privacy\_leakage$, e. Hence, we consider Privacy_NLI as 1- $privacy\_leakage$, denoting the privacy preserved by our method. The higher the metric, the more private information is preserved.
Semantic Relevance. For assessing the preservation of semantic content, we consider ROUGE-1 and ROUGE-Lsum Lin (2004) to compare generated rewrites with the corresponding references.
@@ 4.3 Results and Discussions
We further investigate the rewriting quality w.r.t. each strategy based on human evaluation. We use the T5-Base model trained on the human rewrites and the synthetic data with both strategies, and apply it on the hold-out test set of each strategy. Table 5 shows that the T5-Base model achieves superior performance over the baselines with both strategies. The naturalness of all generated rewrites is on par with that of human rewrites. Both zero-shot Llama-13B models perform better than the best DP method DPNR, which mostly perturbs non-sensitive contents or yields repeated words. The overall results are encouraging for a wide range of applications on edge devices, because our corpus is not huge and T5-Base contains only a few million parameters, which is a few hundred times smaller than Llama-13B, GPT-3.5 turbo and GPT4.
Alignments between Automatic metrics and Human Evaluation. We compare the ranking using Privacy_NLI with the corresponding human judgements in Table 4. T5-NaP2-GPT4 obtains the highest 1-Privacy_NLI of 93.81% in automatic evaluation, matching the highest SPrivacy with 72.00%. The results are aligned well among the rewriting models using the obscuring strategy. However, Privacy_NLI does not rank all rewriting models using deleting in the same manner as humans. To quantify the alignments, we calculate a Spearman’s ranking correlation of 0.70 between Privacy_NLI and SPrivacy among all models to show the effectiveness of Privacy_NLI. The correlation between the models using obscuring reaches even 0.83.
Usefulness of the Synthetic Data. Table 6 shows the result of using synthetic data for training rewriting models. We compare two different strategies: deleting and obscuring. The results shows that the model performs better with the synthetic data for both tasks. In particular, the model preserves more non-personal information compared to human rewrites in the deleting task. With the synthetic data for training the models, the model performance is 7% better than the non-synthetic data model in terms of deleting. The biggest gain of the synthetic data is obtained for improving the privacy protection of the rewriting model using obscuring.
@@ 5 Related Work
The field of controllable text style transfer focuses on modifying specific attributes in texts, such as formality Briakou et al. (2021) and sentiment Li et al. (2018a, 2022) while preserving the core semantic content. The advancement of text rewriting tasks is heavily dependent on the availability of high-quality corpora to assess generation quality. For example,  Rao and Tetreault (2018) collected a large-scale corpus GYAFC for initiating the research of formality style transfer to rewrite formal language. As for our task sensitive to privacy, which demands sophisticated alignment in rewriting utterances, the construction of a specialized corpus for high-quality privacy-sensitive rewrites are crucial.
@@
Another series of work suggested to generate new sentences with less sensitive information Emmery et al. (2018); Xu et al. (2019). Following this approaches, the setting of our work is more general since we use open-domain sensitive personal information as the control signal for rewriting. Moreover, our corpus is flexible in the way that it supports two strategies for rewriting, which is of the interest for the style transfer research community Strengers et al. (2020).
@@ A.1 Question design for human evaluation
A.2 Prompt template for synethetic Data
@@
Figure 2: Prompt template for T5-NaP2.
A.3 Implementation Details
Appendix B Experiments
B.1 Evaluation metrics
@@
  • MAUVE Pillutla et al. (2021): It is widely used as one of the best automatic metrics to quantify the semantic differences between a pair of texts. We adopt it to assess the distribution divergence between model outputs and human written rewrites.
@@
B.2 Impact of DP-SGD.
@@
DP real synth LLM Privacy_NLI MAUVE ROUGE-1 ROUGE-Lsum
False 1300 0 - 0.0810 $\pm$ 0.1077 0.9468 0.6946 0.6924
False 1300 3900 GPT-3 0.0826 $\pm$ 0.0903 0.9558 0.7143 0.7122
False 1300 3900 GPT-4 0.0619$\pm$0.0870 0.9658 0.7301 0.7278
True 1300 0 - 0.0602 $\pm$ 0.0759 0.9584 0.7338 0.7316
True 1300 3900 GPT-3 0.0757 $\pm$ 0.0908 0.9589 0.7368 0.7351
True 1300 3900 GPT-4 0.0703 $\pm$ 0.1135 0.9514 0.7446 0.7428

### v2

@@ 3.2 Evaluation Details
Privacy Leakage. We propose a novel metric, called Privacy_NLI, by using the RoBERTa model trained on the MNLI corpus, to infer to what degree it is possible to infer personal information in personas. As the NLI model classifies a pair of input texts into entailed, contradicted, or neutral, we adopt $P(\text{entailed}|{\bm{x}},{\bm{p}})$ as the score of $privacy\_leakage$, e. Hence, we consider Privacy_NLI as 1- $privacy\_leakage$, denoting the privacy preserved by our method. The higher the metric, the more private information is preserved. To validate the effectiveness of this metric, we incorporated recent NLI models trained with variant MNLI corpus to show the consistency. Besides we also considered Sparse-MAX and Sharp-MAX as soft alignment scoreXu et al. (2020), the detailed comparsion can be found in Appendix. A.7.
Semantic Relevance. For assessing the preservation of semantic content, we consider ROUGE-1 and ROUGE-Lsum Lin (2004) to compare generated rewrites with the corresponding references detailed introduction can be found in Appendix. A.4.
@@ 3.3 Results and Discussions
SNatural LLM-Natural
Human_deleting 95.00% 4.14
Llama2-13B _deleting 87.00% 4.67
T5-NaP2-GPT4 _deleting 95.00 % 4.44
DPNR 19.00% 2.01
Human_obscuring 98.00% 4.37
DP-Prompt 0.00% 1.14
DP-BART 2.00% 1.71
FLAIR-SCRUBBING 0.00% 3.05
Llama2-13B _obscuring 86.00% 4.85
T5-NaP2-GPT4 _obscuring 98.00% 4.36
Table 7: LLM as Naturalness Judge compared with Human preference
@@
We further investigate the rewriting quality w.r.t. each strategy based on human evaluation. We use the T5-Base model trained on the human rewrites and the synthetic data with both strategies, and apply it on the hold-out test set of each strategy. Table 6 shows that the T5-Base model achieves superior performance over the baselines with both strategies. The naturalness of all generated rewrites is on par with that of human rewrites. Both zero-shot Llama2-13B models perform better than the best DP method DPNR, which mostly perturbs non-sensitive contents or yields repeated words. The near-zero SPRIVACY scores observed for these methods stem from the nature of noise injection in embeddings or numeric representations. This results in two extremes: either no change in output due to insufficient perturbation or complete distortion of the generated output. The overall results are encouraging for a wide range of applications on edge devices, because our corpus is not huge and T5-Base contains only a few million parameters, which is a few hundred times smaller than Llama2-13B, GPT-3.5 turbo and GPT4.
Alignments between Automatic metrics and Human Evaluation. We compare the ranking using Privacy_NLI with the corresponding human judgments in Table 5. T5-NaP2-GPT4 obtains the highest 1-Privacy_NLI of 93.81% in automatic evaluation, matching the highest SPrivacy with 72.00%. The results are aligned well among the rewriting models using the obscuring. However, Privacy_NLI does not rank all rewriting models using deleting in the same manner as humans. To quantify the alignments, we calculate a Spearman’s correlation of 0.70 between Privacy_NLI and SPrivacy among all models to demonstrate Privacy_NLI. The correlation between the models using obscuring reaches even 0.83.
Naturalness Assessment via LLM-as-judge We also consider LLM as judge to score the naturalness of generated sentence. Specifically, we reuse the question from human questionnaire about naturalness and convert it to prompt template with score scale from 1-5. The prompt can be found in the Appendix A.9 The results in Table 7 demonstrate a generally strong alignment between LLM-judged naturalness and human preference (SNatural), particularly for high-performing models such as Llama2-13B _deleting, T5-NaP2-GPT4 _deleting, and T5-NaP2-GPT4 _obscuring, where LLM scores ($\geq$ 4.36) closely reflect human-rated naturalness ($\geq$ 95%). However, notable discrepancies emerge for low-performing systems like DP-Prompt and FLAIR-SCRUBBING, where LLMs assign moderately high naturalness scores (e.g., 3.05 for FLAIR-SCRUBBING) despite near-zero human ratings. This indicates that while LLMs can approximate human judgments in many cases, they may overestimate the fluency or coherence of outputs that humans find unnatural, underscoring the need for further calibration of LLM-based evaluation frameworks.
Usefulness of the Synthetic Data. Table 8 shows the result of using synthetic data for training rewriting models. We compare two different strategies: deleting and obscuring. The results shows that the model performs better with the synthetic data for both tasks. In particular, the model preserves more non-personal information compared to human rewrites in the deleting task. With the synthetic data training the models, the model performance is 7% better than the non-synthetic data model in terms of deleting. The biggest gain of the synthetic data is obtained for improving the privacy protection of the rewriting model using obscuring.
@@
NLI score with different backbones
Privacy_NLI Mixed corpus MNLI SPrivacy
DPNR 80.54% 53.69% 25%
DP-Forward 68.87% 48.06% 0%
DP-Prompt 71.85% 82.76% 0%
DP-BART 85.87% 80.72% 1%
FLAIR-SCRUBBING 62.16% 61.49% 0%
T5_zeroshot-deleting 61.33% 49.14% 10%
T5_zeroshot-obscuring 61.33% 19.26% 45%
Llama2-13B_zeroshot-obscuring 81.85% 45.23% 16%
Llama2-13B_zeroshot-deleting 87.04% 76.68% 14%
Llama-paraph-deleting 79.42% 65.91% 31%
Llama-paraph-obscuring 82.69% 52.21% 16%
GPT-3.5-obscuring 90.33% 53.51% 61%
GPT-3.5-deleting 61.33% 49.60% 34%
GPT-4-obscuring 95.36% 56.50% 66%
GPT-4-deleting 89.26% 47.59% 49%
Human Rewrite 97.01% 63.20% 82%
Table 9: Privacy_NLI score with DeBERTa as backbone finetuned with mixed entailment corpus and MNLI
In this section, we present the Privacy_NLI scores obtained using the DeBERTa model as the backbone, fine-tuned with different corpora, including a mixed entailment corpus and MNLI, to evaluate alignment with SPrivacy, which serves as our human evaluation metric. As shown in Table 9, the DeBERTa model achieves varying levels of performance across datasets. Models trained on the mixed corpus, such as DP-BART and deleting, achieve Privacy_NLI scores of 85.87% and 95.36%, respectively, with the former reaching 1% and the latter 66% alignment with SPrivacy. Notably, human rewrites achieve a Privacy_NLI score of 97.01%, with an alignment of 82% with SPrivacy. This comparison underscores the capability of different backbones and training strategies to achieve results close to human-level performance about privacy preservation while maintaining alignment with human evaluation metrics.
@@ 4 Related Work
The field of controllable text style transfer focuses on modifying specific attributes in texts, such as formality Briakou et al. (2021) and sentiment Li et al. (2018a, 2022) while preserving the core semantic content. The advancement of text rewriting tasks is heavily dependent on the availability of high-quality corpora to assess generation quality.  Rao and Tetreault (2018) collected a large-scale corpus GYAFC for initiating the research of formality style transfer to rewrite formal language. As for our task sensitive to privacy, which demands sophisticated alignment in rewriting utterances, the construction of a specialized corpus for high-quality privacy-sensitive rewrites are crucial.
@@
Another series of work suggested to generate new sentences with less sensitive information Emmery et al. (2018); Xu et al. (2019). Recent work has explored prompting LLMs to rewrite sentences containing private information, aiming to obscure sensitive content Emmery et al. (2018); Xu et al. (2019); Staab et al. (2024). However, these approaches often rely on LLMs’ internal knowledge and struggle to align with nuanced human privacy expectations Dou et al. (2024). However in these works, the author does not control how private information to be rewritten in explicit way which may weaken the control for required privacy and human preference. Even advanced techniques like self-disclosure abstraction or adversarial anonymization face challenges in achieving robust, user-aligned privacy and often depend on powerful cloud-based models. In contrast, our work study private rewrite via more diversified free text and supports two rewriting strategies, offering a more flexible and general setting Strengers et al. (2020).
@@ A.1 Question design for human evaluation
A.2 PERSONA-CHAT
The PERSONA-CHAT dataset Zhang et al. (2018) is a crowd-sourced corpus designed to facilitate research in personalized open-domain dialogue systems. Each conversation in the dataset involves two speakers, each assigned a distinct persona comprising 4–5 profile sentences. These personas guide the dialogue, encouraging participants to engage in conversations that reflect their assigned characteristics.
The dataset encompasses:
  • 1,155 unique personas, each with at least 5 profile sentences.
  • 10,907 dialogues totaling over 162,000 utterances.
  • A division into training, validation, and test sets, with 100 personas reserved for validation and another 100 for testing.
This structure promotes the development of dialogue agents capable of maintaining consistent and engaging personalities throughout interactions.
A.3 Prompt template for synethetic Data
@@
Example 1: If I ask you to rewrite [example #1] containing personal information [persona #1] by <deleting/obscuring> private information, you should return [target #1] If I ask you to rewrite [example #1] containing personal information [empty] by <deleting/obscuring> private information, you should return [example #1] Example 2: If I ask you to rewrite [example #2] containing personal information [persona #2] by <deleting/obscuring> private information, you should return [target #2] If I ask you to rewrite [example #2] containing personal information [empty] by <deleting/obscuring> private information, you should return [example #2] Example 3: If I ask you to rewrite [example #3] containing personal information [persona #3] by <deleting/obscuring> private information, you should return [target #3] If I ask you to rewrite [example #3] containing personal information [empty] by <deleting/obscuring> private information, you should return [example #3] Rewrite this sentence, deleting any private information.
Only return the rewritten sentence, nothing else.
Private information present is: [input persona]
Sentence to rewrite: [input utterance]
Figure 2: Prompt template for T5-NaP2
A.4 Experiment Details
A.4.1 Evaluation metrics
@@
A.4.2 Baseline Methods
DPNR. It stands for Differentially Private Neural Representation, which applies Laplace noise to distributed representations of words in order to randomly drop sensitive words or replace sensitive words with non-sensitive ones. We compare the cosine similarity between each word in an input utterance with those in the corresponding persona, and pick the top-$k$ most similar ones.
Llama-paraph. Mattern et al. (2022) points out the limitations of word-level LDP and propose to paraphrase input texts with lower temperature to achieve a sentence-level LDP. We implement this approach by using Llama2-13B.
DP-Prompt.Utpala et al. (2023) utilizes zero-shot prompting and large language model to generate document paraphrasing to prevent author de-anonymization attack which comprise the privacy of text owner with predefined utility constrain.
DP-BART. The method is a privatized text rewriting system incorporates LDP. The system leverages the LPD paradigm to perform model rewriting using BART model to protect input data which tackles same challenge like us.
FLAIR-SCRUBBING. we also adapt the scrubbing method used in  Lukas et al. (2023) as our baseline. We set FLAIR-SCRUBBING as baseline to test if the automatic method can effectively remove private information from sentences.
Zero-Shot LLMs. To compare with the LLMs fine-tuned on our corpus, we apply the same prompts to the same pre-trained LLMs without any training. Specifically, we consider T5-Base, Llama2-13B, GPT-3.5 turbo and GPT4 and apply the prompt template introduced in Sec. 2.3. To distinguish from the fine-tuned models, the T5-Base and Llama2-13B in the zero-shot setting is referred to as T5_zeroshot and Llama2-13B_zeroshot, respectively.
T5-NaP2. By using the same prompts as the zero-shot version, we fine tune T5-Base on the training set of the manually curated corpus, with or without augmenting them with synthetic data. The prompts are similar to those used by zero-shot models detailed in A.3.
A.5 Implementation Details
A.6 Impact of DP-SGD.
@@ A.5 Implementation Details
DP Real Synth LLM Privacy_NLI ROUGE-1 ROUGE-Lsum
False 1300 0 - 0.9190 $\pm$ 0.1077 0.6946 0.6924
False 1300 3900 GPT-3 0.9174 $\pm$ 0.0903 0.7143 0.7122
False 1300 3900 GPT-4 0.9381 $\pm$ 0.0870 0.7301 0.7278
True 1300 0 - 0.9398 $\pm$ 0.0759 0.7338 0.7316
True 1300 3900 GPT-3 0.9243 $\pm$ 0.0908 0.7368 0.7351
True 1300 3900 GPT-4 0.9297 $\pm$ 0.1135 0.7446 0.7428
@@
A.7 Private Information Alignment
Three alignment techniques were evaluated to determine their effectiveness in identifying utterance-persona associations: the RoBERTa MNLI entailment model, Sparse-MAX, and Sharp-MAX. The latter two algorithms, originally proposed for token-level alignment, compute an alignment matrix where each entry represents the probability that a token in an utterance leaks information about a token in a persona. Specifically, for a token $i$ in the utterance and a token $j$ in the persona, the alignment score in row $i$, column $j$ denotes the leakage likelihood.
Since our task requires sentence-level alignment rather than token-wise alignment. We modified algorithms to compute sentence-level alignment probabilities over a sample of 200 utterance-persona pairs. The goal was to correctly identify the persona associated with each utterance based on alignment scores, such that the correct utterance-persona pair receives a high score while unrelated pairs do not.
To make binary alignment decisions, a fixed threshold was applied to the alignment scores. Probabilities below the threshold were interpreted as indicating no alignment, and those above it as indicating alignment. An ideal model would achieve perfect alignment performance, with both precision and recall equal to 1.
Table 12: Private information alignment results.
Family Threshold Recall Precision Min Max Mean Frobenius Norm 1-Norm
RoBERTa Entailment 0.20 0.71 0.25 0.00 0.99 0.04 12.81 8.08
0.25 0.69 0.27 0.00 0.99 0.04 12.81 8.08
0.30 0.69 0.28 0.00 0.99 0.04 12.81 8.08
0.35 0.68 0.29 0.00 0.99 0.04 12.81 8.08
0.40 0.68 0.29 0.00 0.99 0.04 12.81 8.08
0.80 0.57 0.35 0.00 0.99 0.04 12.81 8.08
Sparse-MAX 0.20 1.00 0.01 1.00 1.00 1.00 99.50 99.00
0.25 1.00 0.01 1.00 1.00 1.00 99.50 99.00
0.30 1.00 0.01 1.00 1.00 1.00 99.50 99.00
0.35 1.00 0.01 1.00 1.00 1.00 99.50 99.00
0.40 1.00 0.01 1.00 1.00 1.00 99.50 99.00
0.80 1.00 0.01 1.00 1.00 1.00 99.50 99.00
Sharp-MAX 0.20 1.00 0.01 0.30 0.43 0.35 35.77 38.60
0.25 1.00 0.01 0.30 0.43 0.35 35.77 38.60
0.30 1.00 0.01 0.30 0.43 0.35 35.77 38.60
0.35 0.64 0.01 0.30 0.43 0.35 35.77 38.60
0.40 0.02 0.03 0.30 0.43 0.35 35.77 38.60
0.80 0.00 NaN 0.30 0.43 0.35 35.77 38.60
Our empirical analysis revealed that Sparse-MAX and Sharp-MAX did not generalize well to the sentence-level alignment scenario, as shown in Table. 12. This result is unsurprising given that these algorithms were originally designed for fine-grained, token-level applications. Furthermore, their performance may have been affected by the lack of hyperparameter optimization specific to the sentence-level setting. Due to time constraints, a comprehensive exploration of these configurations was not feasible. Nevertheless, future work may revisit these methods as viable options, contingent on further tuning.
In contrast, the RoBERTa-based MNLI entailment model demonstrated strong alignment performance and required minimal adaptation. Following empirical threshold analysis, a decision boundary of 0.3 was selected for determining alignment. This threshold yielded favorable results and served as the primary alignment mechanism in subsequent experiments.
A.8 Limitation of LDP at Inference Time
Typical scenarios for privacy protection at inference time include i) dataset release; ii) sending queries involving sensitive queries to LLMs hosted on untrusted servers. Local DP can be one possible solution to adding noise locally for individual data releasing and it have more relaxed definition for user input. LDP is designed to make local data pairs indistinguishable and work generally on a sample of instances. The mainstream LDP methods add random noise to local examples to balance privacy and utilities. The collection of modified instances are aggregated to obtain certain statistics for target tasks. The aggregation step is important to mitigate the negative effects of noise for information utility. However, privacy protection at inference time does not allow any aggregation operation among a set of instances and requires finding a tradeoff between utility and privacy for individuals. Thus LDP based text rewriting methods either add too much noise to destroy the utility of information or retain original content involving sensitive information. Our experiments demonstrate the SOTA methods based on LDP empirically and show the promising research direction using our dataset.
A.9 Naturalness Judgment Template
We reuse the questionnaire question of naturalness to form the naturalness template for GPT-4o. We rescale the naturalness from 1-5 where 1 means very unnatural and 5 means perfectly natural. We prompt model to generate a JSON like result to score the rewritten sentence and provide the explanation on it. The detailed template is shown Figure. 3
Utility test for downstream task To further evaluate the utility of rewrites for LLMs, we conducted additional experiments to compare the LLM’s responses generated by original inputs and their rewrites as the downstream task. The original PERSONA-CHAT is designed in a multiple round and chit-chat manner. We locate the input utterance in our datasets in the position of dialogue and compare the generated response with original response which can be formed as the response generation task. Specifically, we feed original texts and their rewrites respectively to the llama3.2-3B-Instruct and compare their responses with candidate responses collected from PersonaChat dataset, from which we sampled our dataset. The candidate set contains both ground-truth and implausible responses written by human.
Generated responses were ranked by calculating the cosine similarity of their embeddings to those of ground-truth responses, using a SentenceBERTReimers and Gurevych (2019). As shown in Table.  14, the response similarity between the original inputs and their rewrites is 35.99%. Both the responses generated from the original inputs and those from the rewrites achieve comparable similarity scores to the ground truth. It is also worth noting that the responses generated from the rewrites are closer to the ground truth than the implausible ones. Therefore, our rewrites achieve similar utility as original inputs.
Generalization of Datasets To further evaluate the generalization ability of our dataset, we consider the rewriting model BART as our backbone and fine-tune the model on our datasets. BART is an efficient model compared to state-of-the-art language models, with only 175 million parameters. As shown in Table 13, the results demonstrate that a generative model with a small parameter size can still effectively adopt human rewrites from our dataset. The Privacy_NLI score is 85.74%, which is competitive with GPT-3.5, and the ROUGE-Lsum score is 57.32%, indicating higher consistency in language compared to Llama-paraph. Considering the gap in parameter size, the results show that our dataset can enable models like BART and T5 to achieve competitive performance in human-like rewriting tasks.
Privacy_NLI ROUGE-1 ROUGE-Lsum
GPT-3.5-deleting 74.29% 69.13% 68.48%
Llama-paraph-deleting 76.42% 56.29% 54.91%
BART-NaP2 85.74% 59.435 57.32%
T5-NaP2-NaP2 93.81% 73.01% 72.78%
Table 13: NaP2 finetuned BART with baseline method
Original Responses with Rewrites Implausible Responses
Ground Truth Responses 82.00% 76.00% 95.00%
Rewrite 34.00% 94.00% 72.00%
Table 14: Response similarity Comparison between varying types of inputs
LLM Naturalness Judgment Prompt You are an expert linguist. Your task is to assess the naturalness of a given sentence — how fluent, human-like, and typical it sounds in everyday language use. Rate the sentence on a scale from 1 to 5: 1 = Very unnatural (full of grammatical mistakes, incoherent) 2 = Mostly unnatural 3 = Somewhat natural (acceptable, but some issues) 4 = Mostly natural (minor issues) 5 = Very natural (fluent, coherent, no errors) Sentence: "Sentence to Assess" Only provide the score and a brief explanation in the following JSON format:
{"score": X, "explanation": "..."}
Figure 3: Prompt used for LLM-based naturalness judgment.